Prove that 2n+1 > (n + 2) · sin(n) for all positive integers n.


1
Expert's answer
2021-08-10T08:51:59-0400
"n>1=>2n>n+1=>2n+1>n+2"

"1\geq\sin n, n>1"

Then


"2n+1>(n+2)\cdot\sin n, n>1"

"n=1, 2(1)+1>(1+2)\cdot\sin (1),"

"3>3\cdot\sin (1)"

"1>\sin (1), True"

Therefore "2n+1>(n+2)\cdot\sin n," for all positive integers "n."